# A Fractal-based Complex Belief Entropy for Uncertainty Measure in Complex Evidence Theory

### v1

@@ V Experiments
Intuitively, the uncertainty of evidence source 1 should be the largest, and that of evidence source 2 should be the smallest. The calculation results of several different entropy of three different evidence sources are shown in Figure 8 and Table IX. From Figure 8 and Table IX, it can be seen that Pal et al.’s entropy, Deng entropy and Zhou et al. s’entropy cannot reflect the change in the amount of information caused by the intersection of focal elements. Cui et al. s’entropy and the proposed FCB entropy can reflect this feature, and the change trend of FCB entropy is more obvious than Cui et al. s’entropy, so FCB entropy is more effective and has a good description ability for uncertain information.
TABLE IX: Results of different entropy in Example V.2.
Entropy
$\mathbb{M}$ Pal et al. [37] Deng [35] Zhou et al. [64] Cui et al. [62] FCB
$\mathbb{M}_{1}$ 2.4598 3.0448 2.6841 3.0448 3.0448
$\mathbb{M}_{2}$ 2.4598 3.0448 2.6841 3.0100 2.6263
$\mathbb{M}_{3}$ 2.4598 3.0448 2.6841 3.0287 2.8508
Figure 8: Line chart of calculation results of different entropy in Example V.2.
Example V.3.
Given a FoD $X=\{{{x}_{1}},{{x}_{2}},\cdots,{{x}_{10}}\}$ and a CBBA $\mathbb{M}$ defined on the FoD is given as follows
$\mathbb{M}({{x}_{1}})=\alpha+0.3i,$
$\mathbb{M}({{X}_{i}})=1-\alpha-0.3i.$
The complex mass assignment in $\mathbb{M}$ varies with the number of elements in ${{X}_{i}}$ ,which is shown in Table X. In addition, the complex mass assignment in $\mathbb{M}$ varies with $\alpha$. Next, set the value of $\alpha$ to 0, 0.1. In the case of taking two different values of a, the results of FCB entropy, Deng entropy [35], Zhou et al.’s entropy [64] changing with the change of $i$ are shown in Figures 9a, 9b.
TABLE X: Results of different entropy in Example V.2.
$i$ $X_{i}$
1 $x_{1}$
2 $x_{1}$, $x_{2}$
3 $x_{1}$, $x_{2}$, $x_{3}$
4 $x_{1}$, $x_{2}$, $x_{3}$, $x_{4}$
5 $x_{1}$, $x_{2}$, $x_{3}$, $x_{4}$, $x_{5}$
6 $x_{1}$, $x_{2}$, $x_{3}$, $x_{4}$, $x_{5}$, $x_{6}$
7 $x_{1}$, $x_{2}$, $x_{3}$, $x_{4}$, $x_{5}$, $x_{6}$, $x_{7}$
8 $x_{1}$, $x_{2}$, $x_{3}$, $x_{4}$, $x_{5}$, $x_{6}$, $x_{7}$, $x_{8}$
9 $x_{1}$, $x_{2}$, $x_{3}$, $x_{4}$, $x_{5}$, $x_{6}$, $x_{7}$, $x_{8}$, $x_{9}$
10 $x_{1}$, $x_{2}$, $x_{3}$, $x_{4}$, $x_{5}$, $x_{6}$, $x_{7}$, $x_{8}$, $x_{9}$, $x_{10}$
From the two figures, it can be seen that when $i=1$, all the mass is allocated to the single element focal element ${{x}_{1}}$. From the logical judgment, ${{x}_{1}}$ is certain, and the uncertainty should be 0. However, only the FCB entropy of the three entropy is 0, which is consistent with objective cognition. As $i$ increases, FCB entropy is basically consistent with Deng entropy. That is, FCB entropy has the performance of measuring uncertainty similar to Deng entropy. And with the increase of $\alpha$, the mass distribution of ${{x}_{1}}$ increases and the uncertainty decreases. The above three kinds of entropy can well reflect this phenomenon. To sum up, FCB entropy has advantages in measuring uncertainty.
(a) When $\alpha=0$, the curves of three entropies changes with $i$.
(b) When $\alpha=0.1$, the curves of three entropies changes with $i$.
Figure 9: Three entropies change with $i$ when $\alpha$ sets different values.
@@ VI Conclusions
As a generalization of D-S evidence theory, complex evidence theory has played an important role in many fields. However, the measurement of uncertainty is still an open issue. Many studies have proposed different measurement methods, such as belief entropy, divergence, etc. As a new subject, fractal theory has attracted many scholars’ interest. In order to link CBBA with probability, a CPBT method is proposed to convert CBBA into probability. However, CPBT only gives the result of conversion without specific process. Therefore, the relationship between CBBA and probability cannot be comprehensively understood. Inspired by fractal theory, this paper proposes a fractal generation process of CPBT. Then, based on the generation process of CPBT, a new basic belief assignment called FCBBA is defined, and on this basis, FCB entropy is proposed to measure the uncertainty of CBBA. The discord and non-specificity are also defined. Then the properties of FCB entropy are analyzed, and several examples are used to verify its effectiveness. In the future, the application of FCB entropy in pattern recognition, evidence fusion and other fields will be analyzed and realized.
@@ Acknowledgments
This research is supported by the National Natural Science Foundation of China (No. 62003280), Chongqing Talents: Exceptional Young Talents Project (No. cstc2022ycjh-bgzxm0070), Natural Science Foundation of Chongqing, China (No. CSTB2022NSCQ-MSX0531), and Chongqing Overseas Scholars Innovation Program (No. cx2022024).

### v2

@@ V-A Discuss from numerical examples
Intuitively, the uncertainty of evidence source 1 should be the largest, and that of evidence source 2 should be the smallest. In order to compare the effect of FCB entropy and the previously proposed entropies on multi-element intersection, the previous entropies are extended to CBBA using Definition 16. The calculation results of several different entropy of three different evidence sources are shown in Figure 6. From Figure 6, it can be seen that Pal et al.’s entropy, Deng entropy and Zhou et al.’s entropy cannot reflect the change in the amount of information caused by the intersection of focal elements. Cui et al.’s entropy and the proposed FCB entropy can reflect this feature, and the change trend of FCB entropy is more obvious than Cui et al.’s entropy, so FCB entropy is more effective in reducing information loss and has a good description ability for uncertain information.
V-B Application in pattern classification
In [60], a pattern classification approach is presented on the basis of entropy. This section employs this approach to evaluate the novel FCB entropy. The method is divided into three stages:
  • The initial stage involves securing the CBBAs for each piece of data in the test set. Compute the differences in the mean and standard deviation between the original training set and the one augmented with test data. According to the definition in [60], the phase angle is represented by the mean value, while the modulus length corresponds to the standard deviation. The initial CBBA is derived as follows:
    $\mathbb{M}\left(A\right)=\left(\frac{1}{{{e}^{\left|\Delta\delta\right|}}}\right){{e}^{i\left(\left|\Delta\mu\right|\right)}},$ (54)
    where $\Delta\delta$ is the change value of the standard deviation, and $\Delta\mu$ is the change value of the mean. If the test set has $N$ data, $N$ sets of initial CBBAs are obtained. Then, the complex mass functions are normalized to ensure that each group satisfies:
    $\sum{\mathbb{M}\left(A\right)}=1.$ (55)
  • The second stage is to determine the CBBA for the training set. Assuming the training set data comprises $n$ distinct classes, $M$ complex-valued masses must be obtained. For a given category, similar to the first stage, calculate the CBBAs for each data point. Subsequently, calculate the sum of differences in FCBE for each data point relative to all others. Entropy signifies the level of confusion; a smaller change in entropy indicates a better fit for the data set. Thus, the complex mass with the minimal entropy change is identified as the optimal mass for the training set.
  • The final stage classifies the test set data. Optimal CBBAs for the $n$ classes are obtained from the second stage. Mirroring the second stage’s approach, a test set data point is selected, and the FCBE change value is determined using the complex mass function for each class. The smaller the entropy change, the more closely the test data aligns with the respective category. In other words, the test data is assigned to the category with the smallest entropy difference.
We utilized two classic datasets from the UCI repository (http://archive.ics.uci.edu/ml/index.php) for pattern classification: Breast Cancer and Glass. The classification outcomes are depicted in Figure LABEL:fig:cl. The x-axis represents the training set proportion, while the y-axis denotes classification accuracy. The test set remains constant, with the training ratio varying from 0.01 to 0.99 across 99 data points. Figure LABEL:fig:cl(a) illustrates the classification accuracy of the new FCB entropy (FCBE), Deng’s complex-valued Deng entropy (CDE) and quantum belief entropy (QBE) on the Breast Cancer dataset. Figure LABEL:fig:cl(b) pertains to the Glass dataset. It is evident that the new FCB entropy outperforms CDE and QBE in terms of classification accuracy on both datasets.
Deng’s experiment has established that complex entropy yields superior classification results compared to real entropy. Consequently, this paper refrains from conducting repetitive experiments. The enhancement is attributed to the fact that real entropy considers only the first element’s information, directly related aspects such as length, density, and quality. In contrast, complex entropy incorporates the second element’s information, indirectly related details. The phase angle encompasses this indirect information. As demonstrated in the experiment, the phase angle retains the mean’s information. Even if the standard deviation is identical, differing means should classify two data points differently. Therefore, complex entropy accounts for the uncertainty of the second element, rendering it more effective than real entropy. Furthermore, in contrast to Deng’s CDE and Wu’s QBE, the FCBE takes into account the influence of event intersections. In other words, the novel FCB entropy harnesses a greater wealth of information concealed within the intersections of distinct focal elements.
TABLE V: The modulus of different diseases’ CBBAs in the fusion evidence.
Evidence fusion Diseases Evaluation results Entropies
Viral fever Malaria Typhoid QB entropy[43] FCB entropy
$\left|\mathbb{M}(\{{{T}_{1}}\})\right|$ $\left|\mathbb{M}(\{{{T}_{2}}\})\right|$ $\left|\mathbb{M}(\{{{T}_{3}}\})\right|$
${\mathbb{M}_{{{D}_{1}}}}\oplus{\mathbb{M}_{{{D}_{2}}}}$ $0.4105$ $0.4105$ $0.1086$ Cannot be determined 2.0791 2.0636
${\mathbb{M}_{{{D}_{1}}}}\oplus{\mathbb{M}_{{{D}_{2}}}}\oplus{\mathbb{M}_{{{D}_{3}}}}$ $0.5248$ $0.3083$ $0.1648$ Viral fever 2.0833 1.7369
${\mathbb{M}_{{{D}_{1}}}}\oplus{\mathbb{M}_{{{D}_{2}}}}\oplus{\mathbb{M}_{{{D}_{3}}}}\oplus{\mathbb{M}_{{{D}_{4}}}}$ $0.6733$ $0.2541$ $0.1096$ Viral fever 1.8243 1.4138
${\mathbb{M}_{{{D}_{1}}}}\oplus{\mathbb{M}_{{{D}_{2}}}}\oplus{\mathbb{M}_{{{D}_{3}}}}\oplus{\mathbb{M}_{{{D}_{4}}}}\oplus{\mathbb{M}_{{{D}_{5}}}}$ $0.8432$ $0.1494$ $0.0766$ Viral frver 1.3716 1.0286
V-C Application in information fusion
We adopt the decision-making method proposed in the information fusion process in [43], and conduct experimental comparisons between QB entropy and FCB entropy. The specific problem description and algorithm are as follows:
Problem statement: Given a frame of discernment $\Theta=\{{{T}_{1}}$, $\ldots$, ${{T}_{i}}$, $\ldots$, ${{T}_{n}}\}$ and $N$ distinct evidence resources $\mathbb{M}=\{{{\mathbb{M}}_{1}}$, $\ldots$, ${{\mathbb{M}}_{j}}$, $\ldots{{\mathbb{M}}_{N}}\}$, the objective is to identify the target from $\Theta$ by integrating the minimal number of evidence sources necessary.
The decision-making process is outlined as follows:
Step 1: For each piece of evidences $j$, fusion results are computed using a complex evidence combination rule:
$Fus(\mathbb{M})={{({{({{\mathbb{M}}_{1}}\oplus{{\mathbb{M}}_{2}})}_{1}}\oplus\cdots\oplus{{\mathbb{M}}_{j}})}_{j-1}}.$ (56)
Step 2: A threshold value $\sigma$ is predefined for decision-making purposes.
If $\left|\mathbb{M}({{T}_{\delta}})\right|<\sigma$, the target remains undetermined, and further fusion is required.
If $\left|\mathbb{M}({{T}_{\delta}})\right|\geq\sigma$, then ${{T}_{\delta}}$ is considered a potential target:
$\delta=\underset{1\leq i\leq n}{\mathop{\arg\max}}\,\{\left|\mathbb{M}(\{{{T}_{i}}\})\right|\},$ (57)
The potential $\text{The potential target}\leftarrow{{T}_{\delta}}.$
Subsequently, FCB entropy is computed using Eq. (28), to obtain ${{\mathbb{E}}_{FCB}}(Fus(\mathbb{M}))$.
A second threshold $\varepsilon$ is set to determine the conclusion of the fusion process.
If ${{\mathbb{E}}_{FCB}}(Fus(\mathbb{M}))\leq\varepsilon$, the judgment is deemed credible, and the fusion process can terminate.
If ${{\mathbb{E}}_{FCB}}(Fus(\mathbb{M}))>\varepsilon$, the judgment remains indeterminate, necessitating further fusion.
Step 3: Repeat Steps 1-2 until the stopping criteria for the fusion process are met.
Figure 8: Comparison of the two entropies in decision-making.
We explore the practicality of FCB entropy through its application in a medical diagnostic context. The example presented is grounded in the complex model, reflecting the intricate nature of real-world data and application scenarios. This application not only demonstrates the versatility of the algorithm but also underscores its suitability for tackling complex diagnostic problems in the medical field. Additionally, we conducted a comparison with QB entropy [43], an emerging measure of uncertainty within the field of CET. This comparison further underscores the efficacy of the novel approach, bringing its advantages into sharper relief. The results are in Table V.
There are conflicts among the five evidence sources. As the fusion process progresses, the result of target recognition becomes more reliable. When ${\mathbb{M}_{{{D}_{1}}}}$, ${\mathbb{M}_{{{D}_{2}}}}$ and ${\mathbb{M}_{{{D}_{3}}}}$ are fused, according to Table V, it can be seen that $\left|\mathbb{M}(\{{{D}_{1}}\})\right|=0.5248$, $\left|\mathbb{M}(\{{{D}_{2}}\})\right|=0.3083$ and $\left|\mathbb{M}(\{{{D}_{3}}\})\right|=0.1648$. In this case, the patient $P$ is most likely to suffer from Viral fever with the belief value of 0.5248, which satisfies the threshold value 0.5. In the meantime, ${{\mathbb{E}}_{FCB}}(Fus(\mathbb{M}))=1.7369$ satisfy the threshold value 2.0. The recognition result under this fusion condition can be considered reliable. Then the fusion process ends and the decision-making problem is solved.
However, when ${\mathbb{M}_{{{D}_{1}}}}$, ${\mathbb{M}_{{{D}_{2}}}}$ and ${\mathbb{M}_{{{D}_{3}}}}$ are fused, ${{E}_{Q}}{{(Fus(\mathbb{M}))}_{\text{avg}}}=2.0833$ don’t satisfy the threshold value 2.0. The recognition result of the fusion is considered to be unreliable and needs to continue the fusion process. That is to say, using FCB entropy can make correct decisions faster and improve efficiency than QB entropy. And in the subsequent fusion process, we can also see that FCB entropy is very sensitive to changes in the amount of information after fusion, indicating the effectiveness of FCB entropy in measuring the uncertainty of CBBA. The specific comparison is shown in Figure 8.
The greatest contribution of FCB entropy is that it fully considers the impact of multi-element focal elements on its subset of focal elements, so it can mine more information from the frame of discernment, making it more advantageous in pattern recognition, information fusion, etc.
@@ VI Conclusions
Complex evidence theory, as an advancement of the D-S evidence theory, has been pivotal in handling uncertainty across numerous domains. Yet, the precise measurement of uncertainty within complex evidence theory continues to be a subject of active research. This paper introduces a novel approach to link complex basic belief assignment with probability through a fractal-inspired complex pignistic belief transformation. Building on this, we define a new metric, Fractal-based belief (FCB) entropy, designed to quantify the uncertainty inherent in complex basic belief assignment. By examining the discord and non-specificity components, FCB entropy offers a nuanced perspective on uncertainty assessment. The properties of FCB entropy are thoroughly analyzed, and its practical utility is validated through various illustrative examples and practical application in pattern classification and information fusion. Looking forward, the application scope of FCB entropy will be further investigated, with the aim of refining decision-making processes and enhancing the understanding of complex systems within complex evidence theory. Moreover, our future research will indeed consider the decision-making capabilities of CBBAs on the complex plane. We plan to explore the geometric and algebraic properties of CBBAs to better understand their decision-making potential. This includes investigating how the complex plane can be utilized to represent and process belief structures that are inherently multidimensional.
@@ Acknowledgments
The authors would like to express the sincere appreciation to the editor and anonymous reviewers for their insightful comments, which greatly improve the quality of this paper. This research is supported by the National Natural Science Foundation of China (No. 62003280), Chongqing Talents: Exceptional Young Talents Project (No. cstc2022ycjh-bgzxm0070), and Chongqing Overseas Scholars Innovation Program (No. cx2022024).